OpenStudy (anonymous):

Find the slope-intercept form for the line satisfying the following conditions. x-intercept 3, y-intercept 2/5 .

OpenStudy (anonymous):

slope intercept form is y=mx + b, correct? you know from the givens that (3,0) and (0, 2/5) are points on the graph. find the slope using these points: (0-2/5)/(3-0) m=(-2/5)/3= 2/15 so now you now your slope (m=2/15) and they *gave* you the y-int (0, 2/5) so you just plug it in y=mx+b y=(2/15)x + (2/5)

OpenStudy (anonymous):

thank u!